A diffraction grating has 1.2 × 104 rulings or lines uniformly spaced over a width w = 25.4mm. This grating is illuminated at normal incidence by yellow light from a sodium vapour lamp of wavelength 589nm. Determine the greatest angle at which an order can occur on either side of the center of the diffraction pattern?


1
Expert's answer
2021-02-17T12:07:27-0500

The grafting spacing d is,

"d= \dfrac{w}{N}" "= \dfrac{25.4\times 10^{-3}}{1.2\times10^4}=21.167 \times 10^{-7}m = 2117nm"

For diffraction pattern on either side of center n "\geq" 1

For "\lambda" = 589nm We have,

"dsin\theta" = n"\lambda"

"\rightarrow \theta = sin^{-1}(\dfrac{n\lambda}{d})"

"\rightarrow sin^{-1}(\dfrac{1\times589\times10^{-9}}{2117\times10^{-9}})"

"\rightarrow sin^{-1}(0.278)" = 16.14"\degree"

∴Ang "\geq" θ≈16.14"\degree"